OpenStudy (anonymous):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 1 3x - 4y = 10 2 -1 1 -2

hartnn (hartnn):

which method do you prefer to solve system of equations ?

hartnn (hartnn):

substitution ? or elimination ?

OpenStudy (anonymous):

which one is easier?

hartnn (hartnn):

in this case, both are equally easy.

hartnn (hartnn):

lets do, elimination. multiply equation 1 by 4 what u get ?

hartnn (hartnn):

don't want to try to solve on your own ??

OpenStudy (anonymous):

sorry one sec.

hartnn (hartnn):

sure, i'll wait.

OpenStudy (anonymous):

8x+12y=4?

hartnn (hartnn):

yes, now multiply equation 2 by 3 what u get ?

OpenStudy (anonymous):

9x-12y=30

hartnn (hartnn):

good. now add the 2 new equation which you got.

hartnn (hartnn):

do u realise that you have eliminated y by doing that ?

OpenStudy (anonymous):

17x=34 do i divide the other side by 17?

hartnn (hartnn):

yup.

hartnn (hartnn):

and u directly get x=.. ?

OpenStudy (anonymous):

thank you. and you get 2

OpenStudy (anonymous):

What is the value of the y variable in the solution to the following system of equations? 5x + 4y = 1 4x + 3y = -1 do i do the same here?

hartnn (hartnn):

here you eliminate x 1) multiply eq 1 by 4 2) multiply eq 2 by 5 3) subtract.

OpenStudy (anonymous):

i am confues. i got -1 for the y variable and thats not one of the choices

hartnn (hartnn):

20 x+16y = 4 right ?

hartnn (hartnn):

20x+15y =-5 right ?

OpenStudy (anonymous):

yes

hartnn (hartnn):

subtracting., 16y-15y = 4-(-5) y=9

OpenStudy (anonymous):

and you get 1y=-1

OpenStudy (anonymous):

my choices are -7 7 9 -9

hartnn (hartnn):

4-(-5) = 4+5 = 9

OpenStudy (anonymous):

oh crap i forgot to add the two. i only subtracted

hartnn (hartnn):

happens.... any more problems ?

hartnn (hartnn):

and are you understanding what is going on ?

OpenStudy (anonymous):

sure. here i have to use the substitution method. 4x + y = 11 x + 2y = 8

hartnn (hartnn):

isolate x from second equation. what u get ?

OpenStudy (anonymous):

is it 2y=8-x

hartnn (hartnn):

but with that , you isolated nothing......not x, nor y.

hartnn (hartnn):

try to keep x on left side.

OpenStudy (anonymous):

oh ok its x=8-2y

hartnn (hartnn):

yes, now substitute that in 1st equation.

OpenStudy (anonymous):

3

hartnn (hartnn):

yes, y=3 good work :)

hartnn (hartnn):

x=. ?

OpenStudy (anonymous):

2

hartnn (hartnn):

correct.

OpenStudy (anonymous):

For the following system, if you isolated x in the first equation to use the Substitution Method, what expression would you substitute into the second equation? -x + 2y = -6 3x + y = 8

OpenStudy (anonymous):

since x is negative do i need to divide the other side by -1?

hartnn (hartnn):

either divide bot side by -1 or i would recommend you to 1)add x on both sides 2)add 6 on both sides.

hartnn (hartnn):

-x + 2y = -6 2y = -6+x 2y +6 = x.